# Open-FinLLMs: Open Multimodal Large Language Models for Financial Applications

### v1

@@ 3.2 FinLLaMA-Instruct: Domain Task Optimization through Instruction Tuning
3.2.1 Financial Instruction Dataset
To optimize FinLLaMA for downstream domain tasks and instruction following ability, we have assembled an extensive instruction-tuning dataset, totaling 573K samples, specifically tailored for financial applications, as shown in Table 3. The dataset is sourced from four open data sources: (1) FLUPE (Xie et al., 2023b), with 123K samples covering key financial NLP tasks; (2) finred (Liu et al., 2023), with approximately 32.67K examples focused on financial report and document comprehension; (3) MathInstruct (Yue et al., 2023), with 262K examples from 13 distinct mathematical rationale datasets; and (4) Sujet-Finance-Instruct-177k [^4] https://huggingface.co/datasets/sujet-ai/Sujet-Finance-Instruct-177k, which integrates data from 18 different financial NLP datasets. To ensure the uniqueness and quality of the data, we performed rigorous deduplication and filtering, addressing potential overlaps between datasets like FLUPE and Sujet-Finance-Instruct-177k. See Appendix C for more details on the instruction tuning data and data processing.
@@ 3.2.1 Financial Instruction Dataset
FinGPT/Hingpt-finred 32.7K 32.7K
TIGER-Lab/MathInstruct 26.2K 262.0K
sujet-ai/Sujet-Finance-Instruct-177k - 177.0K
Total after deduplication 226.3K 573.0K
3.2.2 Instrution-tuning Details
For instruction tuning, we utilize FinLLaMA as the backbone model and conduct training on 8 NVIDIA A100 80GB GPUs for 6 hours. The model is optimized using Qlora (Dettmers et al., 2024) via AutoTrain [^5] https://github.com/huggingface/autotrain-advanced, configured with a block size and model maximum length of 4096. We train the model over 2 epochs with a batch size of 1 and a learning rate of 0.0002. Parameter-efficient tuning is achieved with LoRA settings of $r=64$, $\alpha=128$, and no dropout, using INT4 quantization. All linear modules are targeted, with right-aligned padding. The AdamW optimizer (Loshchilov and Hutter, 2017), coupled with a cosine scheduler, is used for optimization, along with gradient accumulation set to 4.
@@ 3.3 FinLLaVA: Enabling Multimodal Capabilities via Multimodal Instruction Tuning
Stage Dataset Instructions
Multimodal Alignment ALLaVA-4V(Chen et al., 2024b) 468.0K
OCR-VQA(Wang et al., 2023) 79.0K
SynthTabNet(Nassar et al., 2022) 20.0K
UniChart(Masry et al., 2023) 5.0K
ChartQA(Masry et al., 2022) 20.0K
Chart2Text(Obeid and Hoque, 2020) 30.0K
SFT LLaVA-v1.5-mix665k(Liu et al., 2024b) 665.0K
Evol-Instruct*(Chen et al., 2024b) 143.0K
Total 1430.0K
3.3.1 Multimodal Instruction Data
@@
We curate a diverse multi-modal dataset comprising image, tabular, chart, and text data to ensure comprehensive coverage of various data formats, as shown in Table 4. For image data, we utilize three vision instruction datasets: ALLaVA-4V (Chen et al., 2024b), LLaVA-v1.5-mix665k (Liu et al., 2024b), and OCR-VQA (Wang et al., 2023). For chart data, we integrate subsets from multiple sources: UniChart (Masry et al., 2023), with 5k chart image-text pairs; Chart2Text (Obeid and Hoque, 2020), with 30K chart image-text pairs; and ChartQA (Masry et al., 2022), featuring 20K chart images and their QA pairs. We used GPT-4o to evaluate these datasets and filtered out images most relevant to the financial domain. Further details are provided in Appendix D.2. Different from previous work (Bhatia et al., 2024), which focuses solely on image and chart data, we select a subset of the SynthTabNet (Nassar et al., 2022) dataset, consisting of 20K annotated images of data in tabular layouts. Further details are provided in Appendix D.1. Additionally, we included Evol-Instruct (Chen et al., 2024b), a dataset of 143K pure text instructions, to enhance the model’s generalization capabilities and reduce the risk of hallucination.
3.3.2 Multimodal Instruction Finetuning
We utilize CLIP (Radford et al., 2021) as our visual encoder in conjunction with the FinLlaMA language decoder, fine-tuning the model on our multimodal instruction dataset. Our approach follows the training framework established by LLaVA-1.5 (Liu et al., 2024b), implementing a two-stage instruction-tuning process.
Stage 1 Multimodal Alignment: In this initial stage, we aim to align the vision encoder’s output with the language model’s embedding space. During this phase, both the vision encoder and LLM weights remain frozen. The key objective is to train a two-layer MLP projector to bridge the gap between the vision encoder’s features and the LLM’s embedding. For each input, consisting of an image $X_{v}$, instructions $X_{\text{instruct}}$ that may involve single-turn or multi-turn conversations, and the target answer $X_{a}$, the vision encoder processes the image data to generate a vision feature: $Z_{v}=g(X_{v})$. The MLP projector then maps $Z_{v}$ into the embedding space of the language model: $H_{v}=f_{\text{MLP}}(Z_{v};\theta)$, where $\theta$ represents the trainable parameters of the projector. The training objective is to maximize the auto-regressive likelihood: $\sum_{i}^{L}P_{\theta}(X_{i}\mid X_{v},X_{\text{instruct},<i},X_{a,<i})$, where $L$ is the sequence length of the target answer $X_{a}$, and $X_{\text{instruct},<i}$ and $X_{a,<i}$ are the tokens of instructions and answers preceding the current prediction $X_{i}$.
Stage 2 Supervised Fine-tuning: In the second stage, we continue updating the parameters of both the language model and the MLP projector, while keeping the vision encoder’s parameters frozen. We maintain the same autoregressive training objective as the previous stage but apply it to a different dataset. As shown in Table 4, we utilize LLaVA-v1.5-mix665k and Evol-Instruct as our primary training data sources.
@@ 3.3.2 Multimodal Instruction Finetuning
Training Details: In the multimodal alignment stage, we set the global batch size to 128 and the learning rate to $1\times 10^{-3}$. We employ a warm-up ratio of 0.03, where the learning rate linearly increases from 0 to the initial learning rate, followed by a cosine decay schedule. The training uses bf16 and tf32 precisions to leverage tensor core acceleration and ensure numerical stability. Weight decay is set to 0.0 to prevent over-regularization. We set the model’s maximum length to 2048 tokens. We conduct our training on eight NVIDIA HGX H20 80GB GPUs, with the entire process taking approximately 30 hours for one epoch. In the SFT stage, we set the global batch size to 256 and the learning rate to $2\times 10^{-5}$. A warm-up ratio of 0.05 is used, where the learning rate increases linearly to the initial learning rate, and then follows a cosine decay schedule. Training is performed with bf16 and tf32 precisions to balance computational efficiency and model performance. Weight decay remains at 0.0 to maintain model flexibility. For this stage, we increase the model’s maximum length to 8192 tokens, allowing the model to handle longer sequences during instruction fine-tuning. This stage also runs for one epoch.
@@ 4 Experiments
We conduct extensive evaluations on both general and financial tasks. Unlike previous work like FinTral (Bhatia et al., 2024) that only reported performance of the instruction fine-tuned models, our evaluations include: 1) both zero-shot and few-shot performance of FinLLaMA base model, 2) performance of instruction fine-tuned model FinLLaMA-Instruct, 3) trading performance, and 4) multimodal capabilities.
4.1 Performance of Continual Pretrained FinLLaMA
In this section, we evaluate the continual pretrained FinLLaMA model. We reproduce the results of competitor models where possible and report the best publicly available scores or those we reproduced ourselves.
4.1.1 Zero-shot Performance
In zero-shot scenarios, as shown in Table 5, our evaluation uses 19 datasets covering 11 critical financial tasks. A detailed description of evaluation tasks and detailed prompts for each dataset are available in Appendix F.2.1 and Appendix F.3.1.
Table 5: Datasets used for zero-shot evaluation tasks. The dataset with * is newly created in this paper. Abbreviations: SA for sentiment analysis, CS for credit scoring, CA for claim analysis, and QA for question answering. EM Accuracy means exact match accuracy. MCC means Matthews correlation coefficient.
Task Dataset Test Size Metrics License
SA TSA (Cortis et al., 2017) 561 F1 CC BY-NC-SA 4.0
Classification FOMC (Shah et al., 2023a) 496 F1 CC BY-NC 4.0
Classification FinArg-AUC (Chen et al., 2023) 969 F1 CC BY-NC-SA 4.0
Classification MA (Yang et al., 2020) 500 F1 Public
Classification SC (Mariko et al., 2020) 8,630 Entity F1 CC BY 4.0
Math *MC 50 F1 Public
Math KnowledgeMath (Zhao et al., 2023a) 1,000 EM Accuracy MIT License
Math DocMath-Eval (Zhao et al., 2023b) 3,200 EM Accuracy MIT License
CS German (Feng et al., 2024) 1,000 MCC CC BY 4.0
CS Australian (Feng et al., 2024) 690 MCC CC BY 4.0
CS LendingClub (Feng et al., 2024) 2,690 MCC CC0 1.0
FD ccf (Feng et al., 2024) 2,278 MCC DbCL v1.0
FD ccfraud (Feng et al., 2024) 2,097 MCC Public
Distress polish (Feng et al., 2024) 1,736 MCC CC BY 4.0
Distress taiwan (Feng et al., 2024) 1,364 MCC CC BY 4.0
CA ProtoSeguro (Feng et al., 2024) 2,381 MCC Public
CA travelinsurance (Feng et al., 2024) 3,800 MCC ODbL v1.0
QA ConvFinQA (Chen et al., 2022) 1,490 EM Accuracy MIT License
@@ 4.1.1 Zero-shot Performance
Math MC 15.00 19.30 - 18.00
Math KnowledgeMath 2.30 2.70 - 2.50
Math DocMath-Eval 1.80 3.50 - 3.10
@@
Trading Agent FinMem (Average) 46.46 36.63 - 73.46
Overall, Table 6 (and Figure 5 in Appendix F.3.2) demonstrates that FinLLaMA outperforms the baseline models on most financial tasks, highlighting its robustness and versatility in zero-shot settings. It surpasses its backbone model, LLaMA3-8B, on all tasks, highlighting the effectiveness of continual pre-training with large-scale domain-specific data in enhancing financial knowledge. Additionally, FinLLaMA exceeds the performance of BloombergGPT, despite its larger model size of 50B, and also outperforms the current most capable open-source LLM, LLaMA3.1-8B on most tasks.
FinLLaMA shows exceptional performance in sentiment analysis and classification tasks, demonstrating its proficiency in fundamental financial operations. Its improved accuracy in fact-checking, as shown by the FinFact dataset, highlights its ability to comprehend and evaluate financial information, enabling precise judgments on claims. In math problem-solving tasks, our model shows improvement across all datasets (MC, KnowledgeMath, and DocMath-Eval) compared to LLaMA3-8B.
In credit scoring tasks, performance varies significantly across datasets. Upon manual inspection of prediction results, we find that on the German dataset, both our model and LLaMA3-8B predict all cases as one category, while LLaMA3.1-8B predicts the opposite. On the Australian dataset, our model’s superior performance demonstrates the benefits of continuous pre-training, even with anonymized features. In the LendingClub dataset, our model outperforms LLaMA3-8B, though not LLaMA3.1-8B, likely due to its larger scale training data. Additionally, FinLLaMA excels in fraud detection, financial distress identification, and claim analysis, showcasing its robust capabilities across diverse financial tasks.
Table 7: Datasets used for few-shot evaluation tasks. NER means named entity recognition.
Task Dataset Test Size Metrics License
SA FPB (Malo et al., 2014) 970 F1 CC BY-SA 3.0
SA FiQA-SA (Maia et al., 2018) 235 F1 Public
NER NER (Alvarado et al., 2015) 980 Entity F1 CC BY-SA 3.0
4.1.2 Few-shot Performance
For the few-shot evaluation setting, we use four datasets covering three financial NLP tasks, as shown in Table 7. These tasks are aligned with the BloombergGPT evaluation settings to ensure consistency and comparability. Detailed descriptions of the evaluation tasks can be found in Appendix F.3.
Overall, as shown in Figure 6, FinLLaMA consistently outperforms baseline models across a wide range of financial tasks, demonstrating its robustness and versatility in few-shot settings. In the NER task, FinLLaMA achieves a remarkable F1 score of 82.10, significantly surpassing its backbone model LLaMA3-8B (39.18), LLaMA3.1-8B (49.04), and BloombergGPT (60.82). This highlights the substantial improvement in entity recognition due to continual pre-training. For the FPB dataset, FinLLaMA achieves an F1 score of 70.25 in few-shot settings, outperforming LLaMA3-8B (69.65), LLaMA3.1-8B (13.08), and BloombergGPT (51.07). Similarly, in the FiQA-SA dataset, FinLLaMA scored 75.34, surpassing LLaMA3-8B (52.29), LLaMA3.1-8B (65.39), and BloombergGPT (75.05). In the Headlines dataset for classification, FinLLaMA achieves a score of 85.54, outperforming LLaMA3-8B (80.59), LLaMA3.1-8B (59.95), and BloombergGPT (82.20). These results demonstrate FinLLaMA’s strong ability to classify financial texts accurately with minimal examples.
4.1.3 Trading Performance
We further evaluate the trading performance of FinLLaMA using the FinMem agent framework (Yu et al., 2024) on multi-sourced financial data spanning August 15, 2021, to April 25, 2023. The results are presented in Table 8. The FinMem trading task assesses the LLM’s proficiency in single-asset trading, with Cumulative Return and Sharpe Ratio as the key performance metrics. For more tasks and datasets, please refer to Appendix G.
Table 8: Performance comparison of single-asset trading using FinLLaMA vs. baseline LLMs across multiple stocks. Abbreviations: Metrics including CR (Cumulative Return), SR (Sharpe Ratio), NSR (Normalized Sharpe Ratio), AV (Annual Volatility), and MD (Maximum Drawdown). *Note: The $\uparrow$ symbol next to a metric indicates that higher values correspond to better performance, while $\downarrow$ symbol next to a metric indicates that lower values correspond to better performance.
Metrics Asset Buy & Hold 6 LLaMA3.1-8B LLaMA3-8B Palmyra-Fin-70B FinLLaMA
CR (%$\uparrow$ 1 (Hull, 2007) TSLA -16.33 -43.09 -58.44 -0.25 55.73
COIN -5.62 7.56 -3.19 14.31 17.43
GOOG 8.43 16.06 16.31 13.73 10.98
NIO -35.29 -108.42 3.85 -18.90 46.45
Average 8.25 -31.97 -10.37 2.22 32.65
SR $\uparrow$ 2 (Sharpe, 1994) TSLA -0.48 -1.37 -1.89 -0.01 2.45
COIN -0.10 0.15 -0.07 0.31 0.58
GOOG 0.42 0.81 1.01 1.18 0.69
NIO -0.87 -2.80 0.10 -0.50 1.91
Average -0.26 -0.80 -0.21 0.25 1.41
NSR $\uparrow$ 3 TSLA 42.00 27.17 18.50 49.88 90.83
COIN 48.33 52.50 48.83 55.17 59.67
GOOG 57.00 63.50 66.83 69.67 61.50
NIO 35.50 3.33 51.67 41.67 81.83
Average 45.71 36.63 46.46 54.10 73.46
AV (%) $\downarrow$ 4 (Cochrane, 1988) TSLA 70.15 64.50 63.22 69.72 46.54
COIN 110.47 102.41 98.23 96.06 61.81
GOOG 41.34 40.69 33.00 23.85 32.49
NIO 83.31 79.19 77.05 76.80 49.79
Average 76.32 71.70 67.88 66.61 47.66
MD (%) $\downarrow$ 5 (Ang and Chen, 2003) TSLA 55.32 57.22 64.49 47.83 18.83
COIN 60.18 64.47 40.56 60.82 41.42
GOOG 21.19 18.62 20.92 12.36 17.41
NIO 44.98 75.06 36.66 33.50 30.08
Average 45.42 53.84 40.66 38.63 26.94
  • 1
    CR is a key trading performance metric, providing a comprehensive insight into investment performance. It reflects the total change of asset value in percentage over time here.
  • 2
    SR is another core metric for evaluating investment performance and adjusting returns for risk. It is calculated by dividing the portfolio’s average excess return over the risk-free rate by its volatility.
  • 3
    NSR is the normalized SR. $NSR=\frac{SR+3}{6}\times 100\%$
  • 4
    AV is calculated as the Daily Volatility (standard deviation of daily logarithmic returns) multiplied by the square root of the typical number of trading days in a year (252).
  • 5
    MD is a risk assessment metric. It measures the largest drop (percentage of drop-down) in an asset’s value, from its peak to its trough.
  • 6
    Buy & Hold strategy is a passive investment approach commonly used as a baseline strategy, where an investor purchases stocks and holds onto them for an extended period regardless of market fluctuations.
@@ 4.1.3 Trading Performance
As shown in Table 8, FinLLaMA outperforms other LLMs with positive Cumulative Return and Sharpe Ratio metrics, demonstrating profitability in dynamic trading environments. It achieves the highest Sharpe Ratio (SR) of over 1, indicating a superior risk-return balance. Additionally, FinLLaMA maintains investment stability with an Annual Volatility of $47.66\%$ and a Maximum Drawdown of $26.94\%$, both lower than other models. This combination of high Sharpe Ratio and low volatility highlights FinLLaMA’s ability to deliver consistent returns with minimized risk, making it highly reliable for trading strategies. These results highlight the significant impact of continual pre-training in enhancing FinLLaMA’s performance. For detailed results, please refer to Appendix H.
4.2 Performance of Instruction Tuned FinLLaMA-Instruct

### v2

@@ 2.2 FinLLaMA-Instruct: Domain Task Optimization through Instruction Tuning
Financial Instruction Dataset To optimize FinLLaMA for downstream domain tasks and instruction following ability, we have assembled an extensive instruction-tuning dataset, totaling 573K samples, specifically tailored for financial applications, as shown in Table 3. The dataset is sourced from four open data sources: (1) FLUPE (Xie et al., 2023b), with 123K samples covering key financial NLP tasks; (2) finred (Liu et al., 2023), with approximately 32.67K examples focused on financial report and document comprehension; (3) MathInstruct (Yue et al., 2023), with 262K examples from 13 distinct mathematical rationale datasets; and (4) Sujet-Finance-Instruct-177k, which integrates data from 18 different financial NLP datasets. To ensure the uniqueness and quality of the data, we performed rigorous deduplication and filtering, addressing potential overlaps between datasets like FLUPE and Sujet-Finance-Instruct-177k. See Appendix C for more details on the instruction tuning data and data processing.
@@
FinGPT/Hingpt-finred 32.67k 32.67k
TIGER-Lab/MathInstruct 26.2k 262k
sujet-ai/Sujet-Finance-Instruct-177k - 177k
Total after deduplication 226.3k 573k
Instrution-tuning Details. For instruction tuning, we utilize FinLLaMA as the backbone model and conduct training on 8 NVIDIA A100 80GB GPUs for 6 hours. The model is optimized using Qlora (Dettmers et al., 2024) via AutoTrain  [^3] https://huggingface.co/autotrain, configured with a block size and model maximum length of 4096. We train the model over 2 epochs with a batch size of 1 and a learning rate of 0.0002. Parameter-efficient tuning is achieved with LoRA settings of $r=64$, $\alpha=128$, and no dropout, using INT4 quantization. All linear modules are targeted, with right-aligned padding. The AdamW optimizer (Loshchilov and Hutter, ), coupled with a cosine scheduler, is used for optimization, along with gradient accumulation set to 4.
@@ 2.3 FinLLaVA: Enabling Multimodal Capabilities via Multimodal Instruction Tuning
Stage Dataset Source Instructions
Multimodal Alignment ALLaVA-4V (Chen et al., 2024b) 468k
OCR-VQA (Wang et al., 2023) 79k
SynthTabNet (Nassar et al., 2022) 20k
UniChart (Masry et al., 2023) 5k
ChartQA (Masry et al., 2022) 20k
Chart2Text (Obeid and Hoque, 2020) 30k
SFT LLaVA-v1.5-mix665k (Liu et al., 2024b) 665k
Evol-Instruct * (Chen et al., 2024b) 143k
Total 1430k
@@
Multimodal Instruction Data. We curated a multi-modal dataset combining image, tabular, chart, and text data for comprehensive coverage, as shown in Table 4. For image data, we used ALLaVA-4V (Chen et al., 2024b), LLaVA-v1.5-mix665k (Liu et al., 2024b), and OCR-VQA (Wang et al., 2023). For chart data, we integrated subsets from UniChart (Masry et al., 2023), Chart2Text (Obeid and Hoque, 2020), and ChartQA (Masry et al., 2022). We filtered images relevant to the financial domain using GPT-4o. Unlike prior work (Bhatia et al., 2024), we also included 20k annotated images of tabular data from SynthTabNet (Nassar et al., 2022), and 143k text instructions from Evol-Instruct (Chen et al., 2024b) to improve model generalization and reduce hallucinations. Further details are in Appendix D.2 and D.1.
Multimodal Instruction Finetuning. We use CLIP (Radford et al., 2021) as the visual encoder and FinLLaMA as the language decoder, fine-tuning on our multimodal instruction dataset. Our approach follows LLaVA-1.5’s framework (Liu et al., 2024b), using a two-stage instruction-tuning process: (1) Multimodal Alignment: We align the vision encoder’s output with the language model’s embedding space by freezing both model weights and training a two-layer MLP projector to map visual features into the language model’s space. The training objective maximizes the auto-regressive likelihood: $\sum_{i}^{L}P_{\theta}(X_{i}\mid X_{v},X_{\text{instruct},<i},X_{a,<i})$. (2) Supervised Fine-tuning: In this stage, we update both the language model and MLP projector while keeping the vision encoder frozen, applying the same objective to a new dataset, primarily LLaVA-v1.5-mix665k and Evol-Instruct.
@@
Training Details: In the multi-modal alignment stage, we set the global batch size to 128, the learning rate to $1\times 10^{-3}$, with a warm-up ratio of 0.03 and cosine decay. Training uses bf16 and tf32 precision for stability and acceleration. Weight decay is set to 0.0, and the model’s maximum length is 2048 tokens. We train on eight NVIDIA HGX H20 GPUs, completing one epoch in approximately 30 hours. In the SFT stage, we set the global batch size to 256, the learning rate to $2\times 10^{-5}$, with a warm-up ratio of 0.05 and cosine decay. The model’s maximum length is increased to 8192 tokens for longer sequences. Weight decay remains at 0.0, and training runs for one epoch with the same precision settings for efficiency and performance.
@@ 3 Experiments
We conducted broad evaluations on general and financial tasks. Unlike FinTral Bhatia et al. (2024), which only reported instruction-tuned performance, we assess: 1) zero- and few-shot performance of the FinLLaMA base model, 2) FinLLaMA-Instruct performance, 3) trading performance, and 4) multimodal capabilities.
3.1 Performance of FinLLaMA
In this section, we evaluate the continual pretrained FinLLaMA model using 24 datasets spanning 11 critical financial tasks, categorized into financial-NLP tasks (Sentiment Analysis, Classification, QA, and Named Entity Recognition) and financial-specific tasks (Misinformation, Math, Credit Scoring, Fraud Detection, Financial Distress, Claim Analysis, and Decision Making). We compare its performance against competitor models by reproducing their results where possible or reporting the best publicly available scores.
@@ 3.1 Performance of FinLLaMA
Math MC 15.00 19.3 - 18.00
Math KnowledgeMath 2.3 2.7 - 2.5
Math DocMath-Eval 1.8 3.5 - 3.1
@@
Agent Decision Making Single-asset trading 45.50 45.50 43.50 67.67
Zero-shot and Few-shot Performance We evaluate FinLLaMA on both zero-shot and few-shot settings. In the zero-shot setting (Table 8, Appendix F.1), we assess 19 datasets across financial-NLP and financial-specific tasks (Appendix F.4.1, F.4.2). As shown in Table 5 and Figure 5 (Appendix 14), FinLLaMA consistently outperforms baselines, surpassing LLaMA3-8B on all tasks and outperforming BloombergGPT (50B) and LLaMA3.1-8B on most, demonstrating strong zero-shot capabilities from domain-specific continual pre-training. It excels in sentiment analysis, classification, fact-checking (FinFact), and math tasks (MC, KnowledgeMath, DocMath-Eval). In credit scoring, performance varies: on the German dataset, FinLLaMA and LLaMA3-8B predicted one class entirely, while LLaMA3.1-8B predicted the opposite; on the Australian dataset, FinLLaMA performs best; on LendingClub, it beats LLaMA3-8B but not LLaMA3.1-8B. It also shows strong results in fraud detection, financial distress identification, and claim analysis.
In the few-shot setting, we evaluate four datasets across three financial-NLP tasks (Table 9, Appendix F.4.2). As shown in Figure 6, FinLLaMA again outperforms all baselines. In NER, it achieves an F1 of 82.10, surpassing LLaMA3-8B (39.18), LLaMA3.1-8B (49.04), and BloombergGPT (60.82). On FPB, it scores 70.25, outperforming LLaMA3-8B (69.65), LLaMA3.1-8B (13.08), and BloombergGPT (51.07). On FiQA-SA, it achieves 75.34 vs. LLaMA3-8B (52.29), LLaMA3.1-8B (65.39), and BloombergGPT (75.05). On Headlines, it reaches 85.54, outperforming LLaMA3-8B (80.59), LLaMA3.1-8B (59.95), and BloombergGPT (82.20), showing strong classification performance even with minimal examples.
@@
Agent We evaluate FinLLaMA on a financial-specific decision-making task using the FinMem framework (Yu et al., 2024), tested on a single-asset trading dataset from August 15, 2021, to April 25, 2023 (Table 12). This task measures Cumulative Return and Sharpe Ratio as key metrics (Appendix G). As shown in Table 12, FinLLaMA outperforms other LLMs with positive Cumulative Return and Sharpe Ratio, demonstrating profitability in dynamic trading environments. It achieves the highest Sharpe Ratio (above 1), indicating superior risk-return balance. FinLLaMA also maintains stability with an Annual Volatility of 0.4766 and a Maximum Drawdown of 0.2693—both lower than other models. This combination of high Sharpe Ratio and low volatility highlights FinLLaMA’s ability to deliver consistent returns with minimized risk, underscoring the benefits of domain-specific continual pre-training. Detailed results are in Appendix H.
3.2 Performance of FinLLaMA-Instruct